OpenStudy (anonymous):

i have less than 4 minutes to answeSolve the exponential equation using logarithms, rounding the answer to the nearest thousandth (3 decimal places). 2 = 7x+3 r plz help guys plz

OpenStudy (anonymous):

3 mintues 40 secs guys

OpenStudy (anonymous):

whoever makes it gets medal!

OpenStudy (wolf1728):

Solve 2 = 7x+3 by logs?

OpenStudy (anonymous):

yup

OpenStudy (jpsmarinho):

madarahd, stay calm...I think this is not the best practice....We not run for medals I think.... apply ln in both sides

OpenStudy (jpsmarinho):

ln2 = ln(7x+3)

OpenStudy (anonymous):

is it supposed to be 2=7^x+3?

OpenStudy (anonymous):

alright thanks and yes roetee

OpenStudy (jpsmarinho):

ln2 = x*ln7 + ln3 is this Roetee?

OpenStudy (jpsmarinho):

no no

OpenStudy (jpsmarinho):

I saw

OpenStudy (jpsmarinho):

2-3 = 7^x

OpenStudy (anonymous):

ln(-1)=x*ln(7)

OpenStudy (jpsmarinho):

but dont have ln of negative number

OpenStudy (anonymous):

hmmm yeah i think jps is right how did you get negative number

OpenStudy (jpsmarinho):

x = ln(-1)/ln(7) x= ln(-8)

OpenStudy (jpsmarinho):

negative number is not possible

OpenStudy (jpsmarinho):

The equation is correct?

OpenStudy (wolf1728):

2=7^x+3 log (2) = 7 * log(x) + 3 can't do addition with logs

OpenStudy (anonymous):

i guess well see its being grade sometime tommorow probably

OpenStudy (wolf1728):

well we tried :-)

OpenStudy (jpsmarinho):

but you not apply log in 3 wolf1728

OpenStudy (jpsmarinho):

aplied*

OpenStudy (wolf1728):

I didn't take the log of 3 because you can't use logs with addition

OpenStudy (wolf1728):

the other equation -1 = 7^x requires log of a negative number - also impossible.

OpenStudy (anonymous):

ill see but thanks wolf for your explantion it really helped me alot cheers!

OpenStudy (wolf1728):

thanks madarahd of course everyone else helped too